OpenStudy (dls):

What is arccos(cos7)=?

OpenStudy (anonymous):

are we using the definition of \(\arccos x\) whose image is \([0,\pi)\)?

OpenStudy (dls):

yes

OpenStudy (anonymous):

in that case recognize \(2\pi\lt7\lt3\pi\) hence \(\arccos(\cos7)=\arccos(\cos(7-2\pi))=7-2\pi\)

OpenStudy (dls):

3pi-7 too?

OpenStudy (dls):

i guess 3pi will change the angle?

OpenStudy (dls):

@oldrin.bataku ?

OpenStudy (anonymous):

@DLS \(\cos(3\pi-7)=\cos(2\pi+\pi-7)=\cos(\pi-7)=\cos(7-\pi)\ne\cos(7)\)

OpenStudy (dls):

okay thanks!